A gas system has pressure , volume and temperature of 1.308 atm, 6399 mL and 517.3 oC, respectively. How many moles of gas are present? Use three significant figures in your answer and do not include units... numbers only!
Answer #1

We have the relation PV = nRT

Here, P = 1.308 atm, V = 6399 mL = 6.399 L, T = 273 + 517.3 = 790.3 K, R = 0.082 L.atm/K.mol

Using above relation,

n = PV/RT = (1.308 x 6.399)/(0.082 x 790.3) = 0.129 mol

Hence, there will be 0.129 mol of gas present.